OpenStudy (anonymous):

The pattern of numbers below is an arithmetic sequence: 14, 24, 34, 44, 54, ... Which statement describes the recursive function used to generate the sequence? The common difference is 1, so the function is f(n + 1) = f(n) + 1 where f(1) = 14. The common difference is 4, so the function is f(n + 1) = f(n) + 4 where f(1) = 10. The common difference is 10, so the function is f(n + 1) = f(n) + 10 where f(1) = 14. The common difference is 14, so the function is f(n + 1) = f(n) + 14 where f(1) = 10.

OpenStudy (jhannybean):

I'd say find the difference, and that will lead you to your answer.

OpenStudy (jhannybean):

to find the difference, \(d\), just subtract the preceding term by the one before it.\[24 - 14 =~?\]\[43-24 = ~?\]\[44-34=~?\] etc...